# FINITE-DIMENSIONAL GAUSSIAN APPROXIMATION FOR DEEP NEURAL NETWORKS: UNIVERSALITY IN RANDOM WEIGHTS

### v1

@@ 2.2 Error for Gaussian weights to Gaussian
which allows us to use the smallest power $(p-2)/(2p-1)$. Now assuming $n_{1},\ldots,n_{\ell}$ are large enough so the final bound in (2.29) is less than two, we apply the smoothing Lemma 2.11 with $k=2d\leqslant 4$ to find
@@
Finally, plugging the bound in (2.28) for $\ell=L-1$ into (2.27) and using the smoothing lemma (again assuming $n_{1},\ldots,n_{L-1}$ large enough for its application, and to simplify the number of powers) implies
@@ 2.3 Stein’s method and smoothing results
Our implementation of Stein’s method follows from the following theorem, which is by now standard. It can easily be read from (Meckes, 2009, Lemmas 1 and 2), and the fundamental ideas go back to Götze (1991). See also Chatterjee and Meckes (2008), Chen et al. (2011), (Goldstein and Rinott, 1996, Chapter 12), (Nourdin and Peccati, 2012, Chapter 4), and Reinert and Röllin (2009).
@@
Let $Z\in\mathbbm{R}^{\mathfrak{D}}$ be a centered multivariate normal random variable with covariance matrix $\Sigma:=(\sigma_{ij})_{i,j=1}^{\mathfrak{D}}$ and define operator on twice-differential functions $\eta:\mathbbm{R}^{\mathfrak{D}}\to\mathbbm{R}$ by
@@
where we write $\partial_{(\cdot)}$ for partial derivatives with respect to the coordinates given by ()($\cdot$)( ⋅ ). Then for any integrable random vector $X\in\mathbbm{R}^{\mathfrak{D}}$, we have
$\mathsf{d}_{3}(X,Z)\leqslant\sup_{\eta\in\widetilde{\mathcal{F}}_{\mathfrak{D}}}\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\mathcal{A}_{\Sigma}\eta(X)\bigr{]}\bigr{\rvert},$
where ~$\widetilde{\mathcal{F}}_{\mathfrak{D}}$ consists of functions $\eta:\mathbbm{R}^{\mathfrak{D}}\to\mathbbm{R}$ with three derivatives satisfying, for $i,j,k\in\{1,\ldots,\mathfrak{D}\}$,
@@
The next lemma records the outcome of a standard Gaussian smoothing argument to bound the $\mathsf{d}_{1}$ metric in terms of the $\mathsf{d}_{3}$ metric. See (Raič, 2018, Section 4.2).
@@
Let $\zeta\in\mathcal{W}_{\mathfrak{D}}$ and $S\in\mathbbm{R}^{\mathfrak{D}}$ be a vector with i.i.d. standard Gaussian entries. For $\varepsilon>0$, define the regularization
$\zeta_{\varepsilon}(u):=\mathop{{}\mathbbm{E}}\mathopen{}[\zeta(u+\varepsilon S)]=\int\zeta(t)\prod_{i=1}^{\mathfrak{D}}\varphi_{u_{i},\varepsilon^{2}}(t_{i})dt,$
@@
$\frac{\partial}{\partial u_{i}}\zeta_{\varepsilon}(u)$ $=\frac{1}{\varepsilon}\mathop{{}\mathbbm{E}}\mathopen{}[S_{i}\zeta(u+\varepsilon S)],$ (2.30)
$\frac{\partial^{2}}{\partial u_{i}\partial u_{j}}\zeta_{\varepsilon}(u)$ =$\begin{cases}\frac{1}{\varepsilon^{2}}\mathop{{}\mathbbm{E}}\mathopen{}[S_{i}S_{j}\zeta(u+\varepsilon S)],&i\not=j\\ \frac{1}{\varepsilon^{2}}\mathop{{}\mathbbm{E}}\mathopen{}[(S_{i}^{2}-1)\zeta(u+\varepsilon S)],&i=j.\end{cases}$ (2.31)
@@
$\lvert\zeta_{\varepsilon}(u)-\zeta_{\varepsilon}(v)\rvert\leqslant\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\lvert\zeta(u+\varepsilon S)-\zeta(v+\varepsilon S\rvert\bigr{]}\leqslant\|u-v\|_{2},$
@@
$\Bigl{\lvert}\frac{\partial}{\partial u_{i}}\zeta_{\varepsilon}(u)-\frac{\partial}{\partial u_{i}}\zeta_{\varepsilon}(v)\Bigr{\rvert}\leqslant\frac{1}{\varepsilon}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\lvert S_{i}\rvert\bigr{]}\|u-v\|_{2},$
@@
and we can conclude that $(1\wedge\varepsilon^{2})\zeta_{\varepsilon}\in\mathcal{F}_{\mathfrak{D}}$. (Note also these three inequalities follow directly from (Raič, 2018, Lemma 4.6).)
@@
$\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}[\zeta(U)]-\mathop{{}\mathbbm{E}}\mathopen{}[\zeta(V)]\bigr{\rvert}$ $\leqslant\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}[\zeta(U)]-\mathop{{}\mathbbm{E}}\mathopen{}[\zeta_{\varepsilon}(U)]\bigr{\rvert}+\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}[\zeta(V)]-\mathop{{}\mathbbm{E}}\mathopen{}[\zeta_{\varepsilon}(V)]\bigr{\rvert}+\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}[\zeta_{\varepsilon}(U)]-\mathop{{}\mathbbm{E}}\mathopen{}[\zeta_{\varepsilon}(V)]\bigr{\rvert}$
$\leqslant 2\varepsilon\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\|S\|_{2}\bigr{]}+\frac{\mathsf{d}_{3}(U,V)}{1\wedge\varepsilon^{2}}\leqslant 2\sqrt{{\mathfrak{D}}}\varepsilon+\frac{\mathsf{d}_{3}(U,V)}{1\wedge\varepsilon^{2}},$
where the second inequality uses that $\zeta$ is $1$-Lipschitz and that that $(1\wedge\varepsilon^{2})\zeta_{\varepsilon}\in\mathcal{F}_{\mathfrak{D}}$. Taking the supremum over $\zeta\in\mathcal{W}_{\mathfrak{D}}$ implies the same inequality holds with left-hand side replaced by $\mathsf{d}_{1}(U,V)$.
@@
Acknowledgment. Balasubramanian gratefully acknowledges support from National Science Foundation through the grant NSF DMS-2413426.

### v2

@@ 2.2 Error for Gaussian weights to Gaussian
which allows us to use the smallest power (p2)/(2p1)(p-2)/(2p-1). Now assuming $n_{1},\ldots,n_{\ell}$ are large enough so the final bound in (2.29) is less than two, we apply the smoothing Lemma 2.11 with $\mathcal{D}=2d\leqslant 4$ to find
@@
Finally, using the bound in (2.28) for $\ell=L-1$ and using the smoothing lemma (again assuming $n_{1},\ldots,n_{L-1}$ large enough for its application, and to simplify the number of powers) we obtain
@@ 2.3 Stein’s method and smoothing results
Our implementation of Stein’s method follows from the following theorem, which is by now standard. It can easily be read from Meckes (2009, Lemmas 1 and 2), and the fundamental ideas go back to Götze (1991). See also Chatterjee and Meckes (2008), Chen et al. (2011), Goldstein and Rinott (1996, Chapter 12), Nourdin and Peccati (2012, Chapter 4), and Reinert and Röllin (2009).
@@
Let $Z\in\mathbbm{R}^{\mathfrak{D}}$ be a centered multivariate normal random variable with covariance matrix $\Sigma:=(\sigma_{ij})_{i,j=1}^{\mathfrak{D}}$ and define the operator on twice-differential functions $\eta:\mathbbm{R}^{\mathfrak{D}}\to\mathbbm{R}$ by
@@
where we write $\partial_{(\cdot)}$ for partial derivatives with respect to the coordinates given by $(\cdot)$. Then for any integrable random vector $X\in\mathbbm{R}^{\mathfrak{D}}$, we have
$\mathsf{d}_{3}(X,Z)\leqslant\sup_{\eta\in\widetilde{\mathcal{F}}_{\mathfrak{D}}}\bigl\lvert\mathbbm{E}\bigl[\mathcal{A}_{\Sigma}\eta(X)\bigr]\bigr\rvert$,
where $\widetilde{\mathcal{F}}_{\mathfrak{D}}$ consists of functions $\eta:\mathbbm{R}^{\mathfrak{D}}\to\mathbbm{R}$ with three derivatives satisfying, for $i,j,k\in\{1,\ldots,\mathfrak{D}\}$,
@@
The next lemma records the outcome of a standard Gaussian smoothing argument to bound the $\mathsf{d}_{1}$ metric in terms of the $\mathsf{d}_{3}$ metric. See Raič (2018, Section 4.2).
@@
Let $\zeta\in\mathcal{W}_{\mathfrak{D}}$ and $S\in\mathbbm{R}^{\mathfrak{D}}$ be a vector with i.i.d. standard Gaussian entries. For $\varepsilon\in(0,1)$, define the regularization
$\zeta_{\varepsilon}(u):=\mathbbm{E}[\zeta(u+\varepsilon S)]=\int\zeta(t)\prod_{i=1}^{\mathfrak{D}}\varphi_{u_{i},\varepsilon^{2}}(t_{i})dt,$
@@
$\frac{\partial}{\partial u_{i}}\zeta_{\varepsilon}(u)$ $=\frac{1}{\varepsilon}\mathbbm{E}[S_{i}\zeta(u+\varepsilon S)],$ (2.30)
$\frac{\partial^{2}}{\partial u_{i}\partial u_{j}}\zeta_{\varepsilon}(u)$ $=\begin{cases}\frac{1}{\varepsilon^{2}}\mathbbm{E}[S_{i}S_{j}\zeta(u+\varepsilon S)],&i\not=j\\ \frac{1}{\varepsilon^{2}}\mathbbm{E}[(S_{i}^{2}-1)\zeta(u+\varepsilon S)],&i=j.\end{cases}$ (2.31)
@@
$\lvert\zeta_{\varepsilon}(u)-\zeta_{\varepsilon}(v)\rvert\leqslant\mathbbm{E}\bigl[\lvert\zeta(u+\varepsilon S)-\zeta(v+\varepsilon S)\rvert\bigr]\leqslant\|u-v\|_{2},$
@@
$\Bigl\lvert\frac{\partial}{\partial u_{i}}\zeta_{\varepsilon}(u)-\frac{\partial}{\partial u_{i}}\zeta_{\varepsilon}(v)\Bigr\rvert\leqslant\frac{1}{\varepsilon}\mathbbm{E}\bigl[\lvert S_{i}\rvert\bigr]\|u-v\|_{2},$
@@
and we can conclude that $\varepsilon^{2}\zeta_{\varepsilon}\in\mathcal{F}_{\mathfrak{D}}$. (Note also these three inequalities follow directly from Raič (2018, Lemma 4.6).)
@@
[…]
$\leqslant 2\varepsilon\mathbbm{E}\bigl[\|S\|_{2}\bigr]+\frac{\mathsf{d}_{3}(U,V)}{\varepsilon^{2}}\leqslant 2\sqrt{{\mathfrak{D}}}\varepsilon+\frac{\mathsf{d}_{3}(U,V)}{\varepsilon^{2}},$
where the second inequality uses that $\zeta$ is $1$-Lipschitz and that that $\varepsilon^{2}\zeta_{\varepsilon}\in\mathcal{F}_{\mathfrak{D}}$. Taking the supremum over $\zeta\in\mathcal{W}_{\mathfrak{D}}$ implies the same inequality holds with left-hand side replaced by $\mathsf{d}_{1}(U,V)$.
@@
Acknowledgment. We thank the anonymous referees and Associate Editor, as well as the Editor, for their helpful comments, which greatly improved the paper. Balasubramanian gratefully acknowledges support from National Science Foundation through the grant NSF DMS-2413426.